I need you help, I don't know how to change word problem in to augmented matrix form.
Investment analysis. A person has $5,000 to invet, part at 5% and the rest at 10%. How much should be invested at each rate to yield %400 per year? Set up a system of equations and solve using augmented matrix methods.
Thank very much.

Answer by checkley77(12844)

.10x+.05(5,000-x)=400
.10x+250-.05x=400
.05x=400-250
.05x=150
x=150/.05
x=3,000 invested @ 10%
5,000-3,000=2,000 invested @ 5%
Proof:
.10*3,000+.05*2,000=400
300+100=400
400=400